U sen sniddde pearum ta jewdu Dutd anoy apuadde a apap pun sasnty ausfes cm a e ad ssanlx y ureignba 0
Answer #1

           2IBr(g) -------------> I2(g)   + Br2(g)

I         2.2*10^-2                  0             0

C        -2x                           +x            +x

E     2.2*10^-2-2x                 +x          +x

        Kp    = PI2 * PBr2/P^2IBr

         8.5*10^-3   = x*x/(2.2*10^-2-2x)

         8.5*10^-3*(2.2*10^-2-2x) = x^2

            x   = 0.0076

       PI2   = x   = 0.0076atm

       PBr2   = x = 0.0076atm

part-B

The partial pressure of I2 = 0.0076atm

part-C

The partial pressure of Br2 = 0.0076atm